OpenStudy (anonymous):

PLEASE HELP ME I ONLY HAVE 20 MINUTESS OMG A school is having a canned food drive. Each class is challenged to collect 180 cans. If there are c classes in the school, which equation could be used to find the total number of cans, T, the school will collect if each class meets the challenge? A. B. C. D.

OpenStudy (anonymous):



zepdrix (zepdrix):

20 minutes oh noessss :O we must hurryyyyy

OpenStudy (anonymous):

YES I HAVE 25 QUESTIONS :((

OpenStudy (anonymous):

ITAS A FINAL INTERIM ASSESMENT IF I DONT PASS THIS I FAIL MY GRADE :(

zepdrix (zepdrix):

so we want 180 cans `per` classroom. 180 for the first, 180 for the next classroom, 180 for the third, and so on... we have 180 + 180 + 180 + .... a whole bunch of them ya? :O we actually have `c` of them being added together, where c is the number of classrooms!

zepdrix (zepdrix):

I'm trying to explain this without just giving you the answer lol

zepdrix (zepdrix):

So if i have a bunch of stuff being added together, lemme give example:\[\Large\rm 20+20+20+20+20\]is the same as \(\Large\rm 20\cdot5\) 5 twenties, ya?

OpenStudy (anonymous):

so confused

OpenStudy (anonymous):

oh yeas i see

zepdrix (zepdrix):

so if i have a bunch of 180's, \(\Large\rm 180+180+180+...\) I'm going to write that as multiplication, ya?\[\Large\rm 180\cdot c\]A bunch of 180's :O

zepdrix (zepdrix):

Ok i kinda spilled the beans -_- whatever onto the next one

OpenStudy (anonymous):

At Random Stationary, the sales records of 24 employees were examined. Twelve of the employees worked the morning shifts and 12 of them worked the afternoon shifts. The following statistical information was calculated from the number of sales of each employee chosen. Morning Shift Workers Afternoon Shift Workers Mean 32 46 Median 36 33 Mode 20 31 Range 44 62 Based on these samples, what generalization can be made? A. Not enough information is provided to draw any of these conclusions. B. The morning shift workers sold less units than the afternoon shift workers. C. The morning shift workers sold the same number of units as the afternoon shift workers. D. The morning shift workers sold more units than the afternoon shift workers.

zepdrix (zepdrix):

it iddn't copy paste well, morning workers is the first column?

OpenStudy (anonymous):

morning workers 32 36 20 44

zepdrix (zepdrix):

hmm this one i dunno >.< need someone smarter hehe

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which of the following is approximately equal to the expression below? A. 4 B. 5 C. 2 D. 3

OpenStudy (anonymous):

46/3 divided by 43/9

zepdrix (zepdrix):

Mmm do you remember your `keep change flip` thing? :O

zepdrix (zepdrix):

\[\Large\rm \frac{46}{3}\div\frac{43}{9}\quad=\quad\frac{46}{3}\times\frac{9}{43}\]We `keep` the first part the same. We `change` the operation from division to multiplication. We `flip` the last guy. ya? :)

zepdrix (zepdrix):

Multiplication fractions straight across, top with top, bottom with bottom,\[\Large\rm =\frac{46\times9}{3\times43}\]

zepdrix (zepdrix):

And use a calculator or something :)

zepdrix (zepdrix):

enchiladas are soooo good -_- aren't they?

zepdrix (zepdrix):

mmm mexican fud

zepdrix (zepdrix):

yes, treat the fraction bar as division

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Indicate which property is illustrated in Step 1. Step 1 8 + (6 + 11) + 0 = 8 + (11 + 6) + 0 Step 2 = (8 + 11) + (6 + 0) Step 3 = (8 + 11) + 6 Step 4 = 19 + 6 A. associative property of addition B. commutative property of addition C. distributive property D. associative property of multiplication

zepdrix (zepdrix):

ya 3 was right, cool c:

zepdrix (zepdrix):

Commutative property looks like this: a + b = b + a they can commute, move around, same result. Associative property looks like this: (a + b) + c = a + (b + c) they can group up differently, same results though. Distributive property looks like this: a(b + c) = ab + ac we gave the a to each term in the brackets.

zepdrix (zepdrix):

Which one do you think? :O

OpenStudy (anonymous):

associaative?>

zepdrix (zepdrix):

in step 1? Hmm it doesn't look like the brackets moved in step 1. It looks like the 6 and 11 traded places, they commuted, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so distributove

OpenStudy (anonymous):

@skullpatrol